### v1

# The neutrino luminosity and energy spectrum of nova outburst  [^1] Released on March, 1st, 2021

Abstract
The nova outburst can produce a large number of neutrinos, whether it is the nuclear reaction process during the explosion or the shock wave acceleration proton process. We study the low-energy nuclear and thermal neutrino luminosity of novae with CO white dwarf (WD) mass ranging from 0.6 to 1.1 $\rm M_{\odot}$ with different accretion rates $\dot{M}$, core temperatures $(T_{\mathrm{C}})$, and mixing degrees. We find that during the accretion phase, low-energy neutrinos are mainly produced by pp chains and plasma decay, and photon luminosity is greater than low-energy nuclear and thermal neutrino luminosity. During the thermonuclear runaway (TNR) phase, low-energy neutrinos are mainly produced by the CNO cycle and photon-neutrino, and the low-energy nuclear and thermal neutrino luminosity far exceeds the photon luminosity. We find that the more massive the WD, the shorter the cycle time and the higher the low-energy nuclear neutrino luminosity. The higher the accretion rate, the lower the low-energy nuclear neutrino luminosity. If the accretion mixing effect is not taken into account, the outburst interval becomes longer, the low-energy nuclear neutrino luminosity will be increased. And for the cooler nova model $(T_{\mathrm{C}}=1\times 10^{7}\rm K)$, the low-energy nuclear neutrino luminosity will be lower during the accretion phase and higher at the TNR. We also predict the neutrino luminosity and energy spectrum of the upcoming recurrent nova T Coronae Borealis (T CrB). We estimate that the next T CrB outburst has a low-energy nuclear neutrino peak luminosity of $2.7\times 10^{8}\ \rm L_{\nu,\odot}$ and a low-energy nuclear neutrino outburst duration of 88 days. In addition, we predict that the high-energy hadronic neutrino flux produced by T CrB nova can not be observed by the current-generation IceCube.
Classical Novae (251) — Neutrino astronomy (1100) — Hertzsprung Russell diagram (725)
1 Introduction
Neutrinos are considered the most abundant particles in the universe and play a crucial role in the evolution of stars(Pandey, 2024). For example, helium flashes (Catelan et al., 1996), thermal pulses (Shi et al., 2020), and cooling of WD (Fontaine et al., 2001) in low-mass stars, late stage evolution of high-mass stars (Woosley et al., 2002), supernova explosion (Janka, 2017), the cooling of neutron stars (Nomoto and Tsuruta, 1981). In addition, the study of neutrinos has significantly contributed to our comprehension of the fundamental properties and of the Universe. They can probe neutrino interactions at high energies, and explore deviations from the Standard Model of particle physics. For example, measuring the cross section of neutrino and nucleon interaction for energies larger than $\rm 1\ TeV$ (Reno, 2023) and prove that Leptoquarks can change the neutrino cross-section (Klein, 2019). High-energy neutrinos are only produced at the same time as high-energy gamma rays (gamma-ray) through the hadron process, so detecting high-energy neutrinos can also determine the origin of high-energy gamma-ray (Song, 2023). Meanwhile, the flavour composition of high-energy neutrinos can be a probe of neutrino magnetic moments (Popov and Studenikin, 2024). Neutrinos carry invaluable information about the existence (or absence) of energetic protons and shed light on the location of the $\gamma$-ray production region (Sahakyan et al., 2014).
Neutrino detection is rapidly developing as a new astronomical messenger. Super Kamiokande (SK) is a detector used to study neutrino Cherenkov radiation from different sources(Fukuda et al., 2003). It primarily detects and analyzes particle interactions within the energy range of several MeV to a few tens of TeV, such as solar neutrino(Abe et al., 2016) and supernova neutrino(Mori et al., 2022). The IceCube detector, the first of its kind with a gigaton-scale capacity, was primarily designed to observe neutrinos originating from the most violent astrophysical phenomena in our universe (IceCube Collaboration et al., 2006). Such as blazar (Collaboration et al., 2018; Tavecchio, 2018; Oikonomou et al., 2019; Cao et al., 2024), gamma-ray bursts (Biehl et al., 2018), black holes (Ma and Wang, 2024; van Velzen et al., 2024), active galactic (Kurahashi et al., 2022) and tidal disruption events (Winter and Lunardini, 2021). These neutrinos usually have extremely high energies, ranging from TeV to PeV and beyond.
Novae are significant sites for the production of neutrinos (Razzaque et al., 2010). The novae are usually binary systems (Warner, 1995), and the primary star is a white dwarf (WD). According to the type of companion star, it can be divided into two categories: one is the classical nova, the companion star is the main sequence star, and the other is the symbiotic nova, the companion star is the red giant (RG). When a sufficient number of layers of material have accumulated, it eventually triggers a thermonuclear runaway (TNR) explosion on the surface of the WD (Hernanz, 2005), and significantly enhances the luminosity of the WD to approximately $\rm 10^{4-5}\ L_{\odot}$ (Gomez-Gomar et al., 1998; Hellier, 2001; Knigge et al., 2011), such a violent nuclear reaction would produce a large number of electron neutrinos. Meanwhile, the shock wave from the nova outburst accelerates the surrounding particles to produce neutrinos. The shock wave can be categorized into two distinct types: internal shock waves and external shock waves. Internal shock wave can be attributed to the collision between slower-moving ejecta and faster-moving ejecta (Metzger et al., 2015). The fast-moving outflow has the potential to collide with a pre-existing dense wind emanating from the RG star, resulting in the generation of an external shock (Lü et al., 2011; Martin and Dubus, 2013). Symbiotics are therefore more likely to produce meaningful hadronic emission at their forward shocks. However, classical novae can and do produce hadronic emission at internal shocks (Metzger et al., 2015; Ackermann et al., 2014; Cheung et al., 2016; Franckowiak et al., 2018). In the “hadronic” scenario, the particles accelerated by the shock wave undergo interactions with the surrounding matter (proton-proton (pp) or proton-gamma ($\rm p\gamma$) collisions), resulting in the generation of neutral and charged muons. These muons subsequently decay into high-energy $(\rm>1\ GeV)$ neutrinos and gamma rays (Abbasi et al., 2022). The maximum energies of high-energy particles will be contingent upon the duration of the nova, the efficiency of the acceleration mechanism, and the energy losses incurred during cooling (Acciari et al., 2022). The particles are accelerated by external shock waves to produce high-energy neutrinos, which typically occur in symbiotic novae (De Sarkar et al., 2023).
T Coronae Borealis (T CrB) is a classical symbiotic recurrent nova (Iłkiewicz et al., 2023). The primary star of this type of nova is in close proximity to the Chandrasekhar limit, and it must undergo accretion at exceedingly high rates, approximately $1.7\times 10\ \rm M_{\odot}\ yr^{-1}$, in order to rapidly accumulate critical-mass envelopes and consequently outburst with such frequency (Kato, 1991). In addition to T CrB, there are three with red giant donors in the 10 known recurrent novae of the Galactic: V745 Sco, RS Ophiuchus, and V3890 Sgr. The WD mass in T CrB is $1.2\pm 0.2\ \rm M_{\odot}$, the companion red giant has a mass of the red giant of $1.12\pm 0.23\ \rm M_{\odot}$. The average accretion rate $(\dot{M}_{\mathrm{a}})$ during the superactive state is $2\times 10^{-8}\ \rm M_{\odot}\ yr^{-1}$ and a maximum of $4\times 10^{-8}\ \rm M_{\odot}\ yr^{-1}$ (Zamanov et al., 2023). The distance from T CrB to Earth is approximately 914 pc. And many groups predict it will outburst in $2024$ (Schaefer, 2023; Maslennikova et al., 2023). Prior to this, the T CrB is recorded eruptions in 1217.8, 1789.9, 1866.4 and 1946.1, the time span between bursts is approximately 80 years (Schaefer, 2023). Because T CrB is very close to us and is about to outburst, it is necessary to study its neutrino luminosity and energy spectrum.
In this paper, we plan to investigate neutrino luminosity and energy spectrum during nova outburst. Section 2 gives the details of software instruments, input physics, reaction networks and model selection. In section 3 we discuss the low-energy nuclear and thermal neutrino luminosity during nova outburst and the influence of different $M_{\rm{WD}}$, $\dot{M}$ and $T_{\mathrm{C}}$ on low-energy nuclear neutrino luminosity. In section 4 we predict the neutrino energy spectrum when the T CrB next outbursts. The conclusions are given in section 5.
2 Nova MODEL
Classical novae eruptions are the result of unstable hydrogen burning on the surface of either CO WD or ONeMg WD cores, which accrete hydrogen-rich materials from their main sequence (MS) or red giant (RG) companions in low-mass, close binary systems. We employ the open-source stellar evolution code (MESA, version 23.05, Paxton et al. 2011, 2013, 2015, 2018; Paxton 2019; Jermyn et al. 2023) to construct nova models. The MESA is capable of generating state-of-the-art simulations of novae (Denissenkov et al., 2013; Zhu et al., 2021).
In the nova model, the occurrence of mass loss is triggered when the luminosity ($L$) of a star surpasses the super-Eddington luminosity $(L_{\mathrm{Edd}})$. The mass loss rate is
$\dot{M}=-2\eta_{\mathrm{Edd}}\frac{\left(L-L_{\mathrm{Edd}}\right)}{v_{\mathrm{esc}}^{2}},$ (1)
[…]
The neutrino normalization $L_{\nu,\odot}=0.02398\cdot L_{\gamma,\odot}=9.1795\times 10^{31}\mathrm{erg\ s}^{-1}$ is adopted in Farag et al. (2020).
2.1 Neutrinos produced by nuclear reaction and thermal processes
The generation of neutrinos inside stars can be broadly categorized into two distinct processes. One is weak nuclear processes, including $\beta^{\pm}$ decay and $e^{\pm}$ capture (Kato et al., 2020):
$\rm nuclei\ EC\ :\ (Z,A)+e^{-}\longrightarrow(Z-1,A)+\nu_{e}$ (2)
$\rm freep\ EC\ :\ p+e^{-}\longrightarrow n+\nu_{e}$ (3)
$\rm PC\ :\ (Z,A)+e^{+}\longrightarrow(Z+1,A)+\bar{\nu}_{e}$ (4)
$\rm\beta^{+}\ :\ (Z,A)\longrightarrow(Z-1,A)+e^{+}+\nu_{e}$ (5)
$\rm\beta^{-}\ :\ (Z,A)\longrightarrow(Z+1,A)+e^{-}+\bar{\nu}_{e}$ (6)
Such as $\rm p(p,e^{+}\nu)d$, $\rm p(e^{-}p,\nu_{e})d$, $\rm{}^{3}He(p,e^{+}\nu_{e})\alpha$ in pp chains, $\rm{}^{13}N(,e^{+}\nu_{e})^{13}C$, $\rm{}^{15}O(,e^{+}\nu_{e})^{15}N$, $\rm{}^{17}F(,e^{+}\nu_{e})^{17}O$ in CNO cycle. For the nova model, we choose pp-and-cno-extras.net as a network of nuclear reactions during nova outburst. The occurrence of classical novae eruptions can be attributed to the instability in hydrogen burning on the surface of WD cores. Therefore, the most important nuclear reactions during nova eruption are the pp chain and the CNO cycle (Denissenkov et al., 2013).
The other is thermal processes, which are sensitive to temperature and density. Thermal neutrino energy losses are from Itoh et al. (1996), including:
pair annihilation$\rm\left(T>10^{9}\ K\right)$
$e^{+}+e^{-}\longrightarrow\nu+\bar{\nu}$ (7)
photon-neutrino$\rm\left(T<4\times 10^{8}\ K,\rho<10^{5}\ g\ cm^{-3}\right)$
$e^{-}+\gamma\longrightarrow e^{-}+\nu+\bar{\nu}$ (8)
plasma decay$\rm\left(10^{7}<T<10^{8}\ K,10^{4}<\rho<10^{7}\ g\ cm^{-3}\right)$
$\rm\gamma^{*}\longrightarrow\nu+\bar{\nu}$ (9)
bremsstrahlung$\rm\left(10^{8}<\rho<10^{10}\ g\ cm^{-3}\right)$
$e^{-}+(Z,A)\longrightarrow e^{-}+(Z,A)+\nu+\bar{\nu}$ (10)
The MESA nuclear reaction rates are a combination of rates from NACRE (Angulo et al., 1999) and JINA REACLIB Cyburt et al. (2010). The treatment of screening corrections, as proposed by (Chugunov et al., 2007), serves to enhance nuclear reaction rates in dense plasmas. The weak reaction rates utilized in this study are derived from previous works (Langanke and Martínez-Pinedo, 2000; Oda et al., 1994; Fuller et al., 1985).
2.2 Neutrinos produced by $pp$ collisions
T CrB is a recurrent nova, the ambient environments and explosion mechanisms are similar to RS Oph (Evans et al., 2019), when the ejection material encounters the wind material around the red giant, it will cause a shock system (Zamanov et al., 2023).
The interaction between RG wind and the ejecta will produce two shocks, a forward shock propagating into the RG wind, and a reverse shock propagating back into the ejecta (Zheng et al., 2022). According to Aharonian et al. (2022), we also only consider the forward shock for explaining the neutrino emission.
In the early stage, the ejecta expands freely and is almost unaffected by the interstellar medium surrounding the binary. We assume when the distance between the wind and the nova r is much larger than the semi-major axis of the nova $a$, the structure of the wind can be approximately spherical, and the ejecta is assumed to be concentrated at the front of the shock wave with a thickness of $hr_{\mathrm{sh}}$, h is $1/10$ of the shock radius. Meanwhile, we are only interested in the initial day of the T CrB outburst, thus we do not consider the radiative cooling phase after the shock wave deceleration. We assume that the shock radius $(r_{\mathrm{sh}})$ and velocity $(v_{\mathrm{sh}})$ of the T CrB are $\rm 4\times 10^{13}\ cm$ and $\rm 4500\ km/s$ (Zheng et al., 2024), respectively. The T CrB companion has a mass of $1\ \rm M_{\odot}$ (Pavlenko et al., 2020) and a typical mass loss rates $(\dot{M}_{\mathrm{RG}})$ of $5\times 10^{-7}\ \rm M_{\odot}\ \mathrm{yr}^{-1}$ (Ferrarotti and Gail, 2006). The velocity of RG wind $(v_{\mathrm{RG}})$ is $\rm 10\ km/s$.
The accelerated protons in the forward shock will interact with the matter in the RG wind, mainly the collision of protons and protons, and the resulting charged pions and kaons decay produces observable gamma-ray and high-energy neutrinos. The density of the matter in the RG wind can be estimated as:
$n_{\mathrm{RG}}=\frac{\dot{M}_{\mathrm{RG}}}{4\pi r_{\mathrm{sh}}^{2}v_{\mathrm{RG}}m_{\mathrm{p}}}=9.9\times 10^{8}\mathrm{~{}cm}^{-3}\ ,$ (11)
where $m_{\mathrm{p}}$ is the rest mass of the proton.
Neutrino energy spectrum is also sensitive to the maximum accelerated proton energy, and it is to relate the acceleration timescale with the age of the system. If particles are accelerated via diffusive shock acceleration, then the acceleration time should be of order the diffusion time ($D/v_{\mathrm{sh}}^{2}$). For Bohm diffusion, the characteristic time scale of particle acceleration can be estimated as:
$\tau_{\text{acc }}=cR_{\mathrm{L}}/3\xi v^{2}_{\mathrm{sh}}\ \mathrm{~{}s}\ ,$ (12)
where $\xi\leq(v_{sh}/c)^{2}\approx 2\times 10^{-4}$ is acceleration parameter, $R_{\mathrm{L}}=E/eB$ is the Larmor radius of particles in the magnetic field B (in Gauss). According to Bednarek and Pabich (2011) and Marcowith et al. (2018), the proton Larmor radius for a particle energy of 1 GeV can be written as :
$R_{\mathrm{L}}\simeq 3\times 10^{6}E_{\mathrm{GeV}}B_{\mathrm{G}}^{-1}\mathrm{~{}cm}\ .$ (13)
As discussed in Chomiuk et al. (2012), the magnetic field B is
$B=\sqrt{32\pi n_{\mathrm{RG}}k_{\mathrm{B}}T_{\mathrm{RG}}}\ ,$ (14)
where $T_{\mathrm{RG}}\rm$ is the temperature of the RG wind, $k_{B}=1.380649\times 10^{-23}\ \rm J/K$. Based on Stanishev et al. (2004), the RG companion star of T CrB has an effective surface temperature of 3500 K. However, the $T_{\mathrm{RG}}\rm$ produced from the companion star should be less than the surface temperature (Chomiuk et al., 2012). Following De Sarkar et al. (2023) for nova RS Oph, we set $T_{\mathrm{RG}}=1000\rm\ K$. The primary mechanism for hadron energy loss occurs through collisions with the matter present in RG winds. An estimation of the timescale for energy losses due to pion production in $pp$ collisions can be derived from (Diesing et al., 2023):
$\tau_{pp}=\left(\sigma_{pp}k\mathrm{c}n_{\mathrm{RG}}\right)^{-1}\approx 2.2\times 10^{7}(\frac{10^{8}\ \mathrm{cm}^{-3}}{n_{\mathrm{RG}}})\ \mathrm{s}\ ,$ (15)
where $\sigma_{pp}\approx 3\times 10^{-26}\rm\ cm^{2}$ is the total cross section of $pp$ interactions, $k=0.5$ is the inelasticity coefficient in this collision. By comparing the time scale for the energy losses in $pp$ collisions and the characteristic acceleration time scale ($\tau_{\mathrm{pp}}=\tau_{\mathrm{acc}}$), the maximum proton energy can be estimated as :
$E_{p,\max}\approx 1\times 10^{-6}\frac{\xi B\tau_{\mathrm{pp}}v_{\mathrm{sh}}^{2}}{c}\ \rm GeV\ .$ (16)
According to Caprioli (2023), the diffusion length ( $D/v_{\mathrm{sh}}$) of the particles with energy $E_{max}$ can not to exceed the source size $r_{\mathrm{sh}}$, which is similar with Hillas criterion. For our hadronic model, we find that the accelerated proton population can reach a maximum energy of approximately 330 GeV. As described by (Caprioli et al., 2010), the maximum energy is usually the location of the cut-off energies, so we set $E_{0}=330\rm\ GeV$. The results obtained from the nova model discussed in this section will be presented in the subsequent section.
3 Result
Using MESA, we simulate the structure and evolution of 7 nova models including different initial CO WD masses from 0.6 to 1.1 M with a mass interval of 0.1 M, and 1.1 M ONeMg WD. Following Farag et al. (2020), we utilize the neutrino H-R diagram to illustrate the neutrino luminosities throughout the complete life cycles of nova outburst, and discuss how $M_{\mathrm{WD}}$, $\dot{M}$, $T_{\mathrm{C}}$, and mixing affect low-energy nuclear neutrino luminosity of nova.
3.1 neutrino luminosity of nova outburst
Figure 1 shows the nova with 0.9 M CO WD and 1.1 M ONeMg WD singly periodic evolutionary tracks in both the H-R diagram and the neutrino H-R diagram. Several key moments during the nova eruption have been marked in the H-R diagram and neutrino H-R diagram. The $a\to b\to c\to d$ is the explosion curve, the $d\to e\to a$ is the cooling curve. The a $\to$ c clearly illustrates that when the TNR occurs, due to the rapid rise of temperature the intense nuclear reaction produce a large number of neutrinos, and the cross-section of neutrinos interacting with matter is extremely low compared to photons, the low-energy nuclear and thermal neutrino luminosity rises rapidly in a straight line while photon luminosity remains almost constant. It is why in the neutrino H-R diagram, there are two points labeled $a$. When the low-energy nuclear and thermal neutrino luminosity reaches its maximum, it also means that the nuclear reaction rate on the surface of the WD reaches its maximum, the photon luminosity begins to rise, low-energy nuclear and thermal neutrino luminosity would remain constant. The c $\to$ d indicates that when part of the photon’s energy is injected into the shell, the shell begins to expand, the surface temperature of the WD decreases. The particle number density in the nuclear reaction zone decreases, and the nuclear reaction rate begins to decrease, so the low-energy nuclear and thermal neutrino luminosity begins to decrease. The d $\to$ e indicates that photosphere of the WD begins to contract, the effective temperature rises, the nuclear reaction rate remains constant, as does the photon luminosity and low-energy nuclear and thermal neutrino luminosity. Finally, the nova outburst is over. The cooling process of the WD commences as the nuclear reaction concludes, leading to a gradual decline in photon and low-energy nuclear and thermal neutrino luminosity.
Figure 1: The track in the H-R diagram and neutrino H-R diagram for the whole nova multicycle evolution. The left is 0.9 M CO nova model, the right is 1.1 M ONe nova model. Both have center temperature $T_{\mathrm{C}}=3\times 10^{7}\ \rm K$, accrete $25\%$ WD material and $75\%$ solar material at a rate $1\times 10^{-9}\ \mathrm{M}_{\odot}\rm\ yr^{-1}$.
The production of neutrinos can be achieved through both nuclear reactions and thermal processes, as discussed in section 2.1. The left panel of Figure 2 illustrates the ratio of the low-energy nuclear reaction neutrino luminosity to the low-energy thermal process neutrino luminosity, traced along the light curve of the nova eruption. The right panel of Figure 2 shows the ratio of the low-energy nuclear and thermal neutrino luminosity to the photon luminosity, traced along the light curve of the nova eruption. During the accretion phase, the temperature of the accretion layer is less than the Fermi temperature $(7\times 10^{7}\rm\ K)$, the nuclear reactions are weak, neutrinos are mainly produced by thermal processes, the accretion layer of the WD radiates energy mainly through photons. However, when the TNR occurs, the neutrinos produced by nuclear reactions will far exceed those produced by thermal processes. Thus, neutrinos loss becomes the main channel of energy loss at the TNR. In addition, low-energy nuclear and thermal neutrino luminosity far exceeds photon luminosity when the nova outbursts. For example, as our $M_{\mathrm{CO}}=0.9\ \rm M_{\odot}$ nova model, when it outbursts, the peak photon luminosity is $5.5\times 10^{4}\ \rm L_{\odot}$, the peak low-energy nuclear neutrino luminosity is up to $3.6\times 10^{6}\ \rm L_{\gamma,\odot}$.
Since the low-energy nuclear neutrino luminosity is much larger than low-energy thermal neutrino luminosity during nova outbursts, we only discuss the low-energy nuclear neutrino luminosity during nova outburst later.
Figure 2: The low-energy neutrino luminosity produced by the nuclear reaction and thermal processes (the left panel). The colorful solid lines are the ratio of the low-energy nuclear reaction neutrino luminosity to the low-energy thermal neutrino luminosity plotted along the light curve of the nova eruption. Gray curves indicate where nuclear reaction neutrinos dominate, green curves where thermal neutrinos dominate, and blue and red curves where the reaction and thermal neutrino luminosities are within a factor of 10. The right panel shows the ratio of the low-energy nuclear and thermal neutrino luminosity to the photon luminosity during the whole nova multicycle evolution. The colorful solid lines give the ratio of the low-energy nuclear and thermal neutrino luminosity to the photon luminosity in a light curve of the nova eruption. The nova model has a 0.9 $\rm M_{\odot}$ CO WD, center temperature $T_{\mathrm{C}}=3\times 10^{7}\ \rm K$, accrete $25\%$ WD material and $75\%$ solar material at a rate $1\times 10^{-9}\ \mathrm{M}_{\odot}\rm\ yr^{-1}$.
Figure 3 shows the the energy yields of the four thermal neutrino processes, and the nuclear reaction rates of the nine neutrinos producing nuclear reactions in the pp chain and the CNO cycle. The left panel shows that when the pre-TNR accreted mass is determined, for the nuclear process, the proton-proton chain plays a crucial role during the primary accretion phase, especially $pp$ reaction: $\rm p+p\to d+e^{+}+\nu$ (Starrfield et al., 2016) and $\rm{}^{7}Be$ electron capture :$\rm{}^{7}Be(e^{-},\nu)^{7}Li$. As accretion progresses, when the temperature of layers is about $7\times 10^{7}\ \rm K$, the degeneracy is no longer important and the layers begin to expand, the temperature begins to rise rapidly, nuclear reaction neutrinos are dominated by CNO cycle, and the TNR occurs. Neutrinos are produced mainly by four nuclear reactions: $\rm{}^{13}N(e^{+},\nu_{e})^{13}C$, $\rm{}^{15}O(e^{+},\nu_{e})^{15}N$, $\rm{}^{17}F(e^{+},\nu_{e})^{17}O$, $\rm{}^{18}F(e^{+},\nu_{e})^{18}O$. For thermal process, during the accretion phase, the WD has a high density in both its interior and layers, the neutrino production is dominated by plasma decay. Once the TNR occurs, neutrino production will be dominated by photon-neutrino because the temperature of layers is in the temperature range $\rm(T<4\times 10^{8}\ K)$ that it dominates, as shown in the right panel.
Figure 3: The nuclear reactions rates of the nine neutrinos producing nuclear reactions in the pp chain and the CNO cycle during the whole nova multicycle evolution (the left panel). The right panel shows the energy yield of four thermal processes during the whole nova multicycle evolution. The blue line is plasma decay, yellow line is bremsstrahlung, green line is photon-neutrino, cyan line is electro pair annihilation. The nova model is similar to Figure 2.
3.2 Neutrino luminosity of novae with different $M_{\mathrm{WD}}$, accretion rate $\dot{M}$, $T_{\mathrm{C}}$, and mixing
The $M_{\mathrm{WD}}$ of nova can affect temperatures in the nuclear burning region. The left panel of Figure 4 shows the neutrino luminosity curves of the range of CO $M_{\mathrm{WD}}$ is $0.6\ \rm M_{\odot}$ to $1.1\ \rm M_{\odot}$. The right panel shows the average peak low-energy nuclear neutrino luminosity of 10 outbursts. It is clearly illustrates that the peak luminosity of neutrinos increases as the mass of the WD increases. The larger the mass and smaller the radius of a WD, the greater the gravitational acceleration. This makes it harder for hydrogen to penetrate the degenerate WD core. That means a more degenerate hydrogen-rich envelope, higher temperatures, and, of course, a more violent explosion. As discussed above, nuclear neutrino is mainly produced by CNO cycle, CNO cycle is highly sensitive to template ($\varepsilon_{\rm CNO}\propto T^{17}$). So the more massive the $M_{\mathrm{WD}}$, the higher the low-energy nuclear neutrino luminosity during the TNR.
Figure 4: Neutrino luminosity curves with different CO nova mass models. Left: The neutrino luminosity of the burst sequence during 0-400000 year with different CO nova mass models, i.e., 0.6 $\rm M_{\odot}$ (cyan), 0.7 $\rm M_{\odot}$ (grey), 0.8 $\rm M_{\odot}$ (red), 0.9 $\rm M_{\odot}$ (green), 1.0 $\rm M_{\odot}$ (blue), and 1.1 $\rm M_{\odot}$ (yellow). Right: Average light curves. Both nova models have center temperature $T_{\mathrm{C}}=3\times 10^{7}\ \rm K$, accrete $25\%$ WD material and $75\%$ solar material at a rate $\dot{M}=1\times 10^{-9}\ \mathrm{M}_{\odot}\rm\ yr^{-1}$.
The left panel of Figure 5 shows the neutrino luminosity curves of nova models with different $\dot{M}$. The right panel shows the average peak low-energy nuclear neutrino luminosity of 10 outbursts. It is clearly show that the higher the accretion rate $\dot{M}$, the smaller the outburst interval $\Delta t_{\mathrm{rec}}$ and $L_{\nu}$. TNR is produced when the hydrogen-rich material is heated to a high enough temperature $\rm(7\times 10^{7}\ K)$ for the hydrogen to burn through the CNO cycle. The high accretion rate provides heat at a high rate (Prialnik et al., 1982), thereby inducing an acceleration in nova eruptions. However, owing to the reduced mass of accretion, the lower TNR peak temperature eventually gives rise to a diminished low-energy nuclear neutrino luminosity at higher accretion rates compared to lower ones. When the accretion rate is low, the accretion layer temperature rises slowly and the outburst interval is larger, which means that the accretion layer accrets more material, the outburst intensity is higher, and the low-energy nuclear neutrino luminosity produced during the outburst is higher.
Figure 5: Neutrino luminosity curves with different $\dot{M}$. Left: The neutrino luminosity of the burst sequence during 0-300000 year with different $\dot{M}$, i.e., $1\times 10^{-9}\ \mathrm{M}_{\odot}\rm\ yr^{-1}$ (cyan), $2\times 10^{-9}\ \mathrm{M}_{\odot}\rm\ yr^{-1}$ (grey), $3\times 10^{-9}\ \mathrm{M}_{\odot}\rm\ yr^{-1}$ (red), $4\times 10^{-9}\ \mathrm{M}_{\odot}\rm\ yr^{-1}$ (green). Right: Average light curves. Both nova models have center temperature $T_{\mathrm{C}}=3\times 10^{7}\ \rm K$, accrete $25\%$ WD material and $75\%$ solar material at a 0.9 $\rm M_{\odot}$ CO WD.
The left panel of Figure 6 displays the neutrino luminosity curves of a hot model $(T_{\mathrm{C}}=3\times 10^{7})$ and cold models $(T_{\mathrm{C}}=2\times 10^{7}\ \mathrm{K}\ \mathrm{and}\ T_{\mathrm{C}}=1\times 10^{7}\ \rm K)$ and with $0.9M_{\odot}$ WD. The right panel shows the average peak low-energy nuclear neutrino luminosity of 10 outbursts. When the initial core temperature of the nova model WD is lower, most of the gravitational energy released by the accretion compressed material will be directed to the core, which will delay the time to reach the TNR. Of course, it means the envelope will accrete more H-rich materials, the outbursts are more intense and the low-energy nuclear neutrino luminosity is higher. However, due to the lower temperature of the core and accreted layer, the nuclear reaction rate during accretion is also lower, resulting in a low neutrino luminosity during the accrete phase. In the hot WD $(T_{\mathrm{C}}=3\times 10^{7}\ \rm K)$, convection zones will appear in the outer core layer. The accreted hydrogen rapidly mixes in the convective region and continues to diffuse to the core, leading to an early outburst (Schwartzman et al., 1994). However, because the accretion mass is less and the temperature is lower, the low-energy nuclear neutrino luminosity produced by the outburst is lower.
Figure 6: Neutrino luminosity curves with different $T_{\mathrm{C}}$. Left: The neutrino luminosity of the burst sequence during 0-50000 year with different $T_{\mathrm{C}}$, i.e., $1\times 10^{7}\rm\ K$ (red), $2\times 10^{7}\rm\ K$ (grey), $3\times 10^{7}\rm\ K$ (cyan). Right: Average light curves. Both nova models have accrete rate $1\times 10^{-8}\ \mathrm{M}_{\odot}\rm\ yr^{-1}$, accrete $25\%$ WD material and $75\%$ solar material at a 0.9 $\rm M_{\odot}$ CO WD.
The left panel of Figure 7 shows the neutrino luminosity curves of nova models when mixing is considered or not. The right panel shows the average peak low-energy nuclear neutrino luminosity of 10 outbursts. If the mixing of accretion material and core CO-enriched material is considered during the accretion stage, it means an increase in the metallicity of the accretion material. On the one hand, increasing the metallicity of the accreting material results in an increase in the opacity. The increased opacity not only allows more heat from the compression of accretion material to be trapped in the accretion layer, but also more heat from nuclear burning to be trapped in the nuclear reaction zone so that the temperature in the nuclear burning region increases faster and the shorter the time to reach TNR than ignore mixing nova model. Therefore, the total amount of accreted and ejected mass is less (Starrfield et al., 1998). On the other hand, mixing of CO-enriched materials can also outburst earlier because it acts as a catalyst for the CNO cycle (H burning). Models that do not take into account mixing require more accretion matter because the temperature rises more slowly in the nuclear burning region. Thus, compared to nova models that consider mixing, the nova model that does not consider mixing has a denser accretion layer, more intense outbursts, and higher low-energy nuclear neutrino luminosity.
Figure 7: Neutrino luminosity curves with different mixing. Left: The neutrino luminosity of the burst sequence during 0-15000 year when accrete $25\%$ WD material and $75\%$ solar material and not consider any mixing. Right: Average light curves. Both nova models have center temperature $T_{\mathrm{C}}=3\times 10^{7}\ \rm K$, 0.8 $\rm M_{\odot}$ CO WD at a rate $\dot{M}=1\times 10^{-8}\ \mathrm{M}_{\odot}\rm\ yr^{-1}$.
3.3 neutrino luminosity of T CrB nova
Considering $M_{\rm WD}$ of T CrB is 1.1-1.3 $\rm M_{\odot}$ and the time span between outbursts is approximately 80 years, the average accretion rate $\dot{M}_{a}\approx 2\times 10^{-8}\ \mathrm{M}_{\odot}\rm yr^{-1}$ (Zamanov et al., 2023), so we use $M_{\mathrm{WD}}=1.3\ \rm M_{\odot}$, $\dot{M}=1.72\times 10^{-8}\ \mathrm{M}_{\odot}\rm\ yr^{-1}$ as our T CrB nova model. The left panel of Figure 8 simulates the low-energy nuclear neutrino luminosity of the first two bursts of the T CrB and the next one shortly after, with peak low-energy nuclear neutrino luminosity of $1.06\times 10^{8}\ L_{\nu,\odot}$, $2.05\times 10^{8}\ L_{\nu,\odot}$, and $2.70\times 10^{8}\ L_{\nu,\odot}$, respectively. In addition, According to the definition of outburst duration in section 2, it can be predicted that the next outburst of T CrB will last about 88 days, as shown in the right panel of Figure 8.
Figure 8: The neutrino luminosity curves of T CrB nova. The left panel is first two bursts of the T CrB nova and the next one shortly after, the right panel is predicted duration of the next outburst of T CrB.
3.4 neutrino energy spectrum
In this section, we consider the neutrino spectrum is produced in two ways by novae, one by nuclear reactions inside novae and the other by hadronic processes outside novae.
3.4.1 nuclear reaction
Considering that the scene of nova outburst is similar to that of supernovae explosion. The CNO cycle that dominates nova outbursts and the fast electron-neutrino bursts $(\mathrm{p}+\mathrm{e}^{-}\longrightarrow\mathrm{n}+v_{\mathrm{e}})$ in supernovae are both beta decay, so we think that the neutrino spectrum produced by the nova nuclear reaction is similar to the supernova neutrino spectrum. Following the Baxter et al. (2022), the initial neutrino energy spectrum has the form:
$\Phi\left(E_{\nu}\right)=\frac{L_{\nu}}{4\pi D^{2}_{\mathrm{TE}}\left\langle E_{\nu}\right\rangle}\frac{(\alpha+1)^{\alpha+1}}{\left\langle E_{\nu}\right\rangle\Gamma(\alpha+1)}\left(\frac{E_{\nu}}{\left\langle E_{\nu}\right\rangle}\right)^{\alpha}\exp\left(-\frac{(\alpha+1)E_{\nu}}{\left\langle E_{\nu}\right\rangle}\right),$ (17)
where $D^{2}_{\mathrm{TE}}$ is the distance between T CrB and Earth, ${\left\langle E_{\nu}\right\rangle}$ is mean energy. The $\alpha$ is spectral shape parameter.
As discussed in section 3.3, the low-energy nuclear neutrino luminosity of the next outburst of T CrB is $2.70\times 10^{8}\ L_{\nu,\odot}$. The CNO cycle produces neutrinos with mean energies of about 2 MeV (Borexino Collaboration et al., 2020). The $\alpha$ is 2.5 (Keil et al., 2003). The electron neutirno spectrum is shown in Figure 9. The model predicts the electron neutrino flux from nuclear reaction of T CrB outburst that can not exceeds the sensitivity limit of Super-Kamiokande.
Figure 9: The estimated total electron neutrino flux reaching the Earth from T CrB (orange line). The Super-Kamiokande’s upper limit on electron neutrino flux (Suzuki, 2019) is blue line.
3.4.2 shock speed
Protons accelerated by the shock of the nova interact with ambient particles in RG wind can produce high-energy neutrinos. The flux of muon neutrinos reaching Earth from the T CrB is calculated using the semi-analytical formulation developed in Kelner et al. (2006).
$\Phi_{\nu}\left(E_{\nu}\right)=\frac{\mathrm{c}n_{\mathrm{RG}}}{4\pi D^{2}_{\mathrm{TE}}}\int_{0}^{1}\sigma_{\text{inel }}\left(E_{\nu}/x\right)J_{p}\left(E_{\nu}/x\right)F_{\nu}\left(x,E_{\nu}/x\right)\frac{dx}{x},$ (18)
where c is light speed. $F_{\nu}\left(x,E_{\nu}/x\right)$ represents a neutrino spectrum produced by a single energy proton, $x$ is $E_{\nu}/E_{p}$. $\sigma_{\mathrm{inel}}$ is total cross section of $pp$ interactions
$\sigma_{\rm inel}(E_{\mathrm{p}})=34.3+1.88L+0.25L^{2}\ mb\ ,$ (19)
where $L=\mathrm{ln}(E_{\mathrm{p}}/1\ \mathrm{TeV})$. $J_{\mathrm{p}}\left(E_{\nu}/x\right)$ represents the distribution of protons
$J_{\mathrm{p}}\left(E_{\mathrm{p}}\right)=\frac{A}{E_{p}^{\alpha}}\exp\left[-\left(\frac{E_{\mathrm{p}}}{E_{0}}\right)^{\beta}\right]\ ,$ (20)
$\alpha$ is power-law index which we select 2.2, $\beta=1$ (Kelner et al., 2006). A is determined based on the given condition
[…]
where the minimum proton energy is the the rest mass energy of proton ($E_{\mathrm{min}}\approx 1\rm\ GeV$). The $E_{\rm{dens,T\ CrB}}$ is the energy density of protons around T CrB, it is determined by the kinetic energy of the shock wave (Acciari et al., 2022):
$E_{k}=0.5M_{ej}v_{sh}^{2}=2.0\times 10^{44}\left(\frac{M_{ej}}{10^{-6}M_{\odot}}\right)\left(\frac{v_{sh}}{4500\mathrm{~{}km}\mathrm{~{}s}^{-1}}\right)^{2}\mathrm{erg}\ ,$ (22)
where $M_{\mathrm{ej}}$ is the total ejection mass during the eruption. Based on our T CrB nova model, $M_{\mathrm{ej}}=1.36\times 10^{-6}\ \rm M_{\odot}$. According to Caprioli and Spitkovsky (2014), the $10\%$ of shock kinetic energy can be converted to the energy of high-energy protons. For T CrB nova, the energy released into accelerated protons ($E_{\rm p,nova}$) is about $2.72\times 10^{43}\ \rm erg$. Meanwhile, the energy density of protons surrounding the T CrB can be determined by dividing the total energy by the volume of the corresponding region:
$E_{\rm{dens,T\ CrB}}=\frac{3E_{p,\text{ nova }}}{4\pi r_{\text{sh}}^{3}}\ .$ (23)
Finally, we can calculate that the proton energy density near the shock wave is $102\ \rm erg/cm^{3}$, which is much greater than that the average energy density in CRs in the Milky Way ( $\sim\ 1\ \rm erg/cm^{3}$).
To validate the credibility of our hadron model, we conduct an estimation on the gamma rays emitted during the RS Oph nova event and compare our results with those presented in De Sarkar et al. (2023). Protons accelerated by the shock of the nova interacting with ambient particles in RG wind also can produce high-energy gamma-ray, $pp$ interactions resulting in the generation of neutral muons, these muons subsequently decay into high-energy gamma-ray ($\pi^{0}\to 2\gamma$). The fitting formula of gamma-ray flux produced by $pp$ interaction has the same form as equation (19):
$\Phi_{\gamma}\left(E_{\gamma}\right)=\frac{\mathrm{c}n_{\mathrm{RG}}}{4\pi D^{2}_{\rm RE}}\int_{0}^{1}\sigma_{\text{inel }}\left(E_{\gamma}/x\right)J_{p}\left(E_{\gamma}/x\right)F_{\gamma}\left(x,E_{\gamma}/x\right)\frac{dx}{x}\ ,$ (24)
here $D_{\rm RE}$ is the distance between RS Oph and Earth, $x$ is $E_{\gamma}/E_{p}$. The key parameters for calculating the gamma-ray emissions of RS Oph include : $v_{\mathrm{sh}}=\rm 4500\ km/s$, $r_{\mathrm{sh}}=\rm 4\times 10^{13}\ cm$, $v_{\mathrm{RG}}=\rm 10\ km/s$, $\dot{M}_{\mathrm{RG}}=5\times 10^{-7}\ \rm M_{\odot}\ \mathrm{yr}^{-1}$, $\tau_{pp}\approx 2.2\times 10^{6}\ \mathrm{s}$, $B=0.11\ \rm G$, $\alpha=2.2$, $E_{0}=400\ \rm GeV$. The gamma-ray flux from RS Oph to Earth is depicted in Figure 10. Our analysis reveals that the MAGIC and H.E.S.S. gamma-ray data points can be consistently explained by gamma-ray generated within our hadronic model.
Figure 10: The estimated gamma-ray flux produced by by the $pp$ interaction from the RS Oph outburst. The orange line is the gamma-ray flux produced by the $pp$ interaction from the RS Oph outburst predicted by De Sarkar et al. (2023). The blue line is the total muonic neutrino flux from the RS Oph outburst by our hadronic model. H.E.S.S. and MAGIC data points are taken from Aharonian et al. (2022), and Acciari et al. (2022), respectively.
The muon neutrino flux after considering neutrino oscillations is illustrated in Figure 11. Neutrino oscillations that its flavor states (electron neutrinos, muon neutrinos, and tau neutrinos) change during propagation. That is, the muon neutrino flux reaching Earth is 1/3 of the total neutrino flux produced by the $pp$ interactions at the T CrB. The blue line is the neutrino flux from the RS Oph outburst predicted by De Sarkar et al. (2023), we reproduce their result and compare it with our predictions of the neutrino flux of the T CrB. The neutrino spectrum generated by the T CrB outburst exhibits similarities to that of RS Oph, as predicted by Gagliardini et al. (2024). However, due to its closer proximity and higher maximum energy of proton production during the outburst, the resulting neutrino flux reaching Earth is amplified. As shown in Figure 11, it is unlikely that the current-generation IceCube will observe it.
Figure 11: The estimated total muonic neutrino flux reaching the Earth from T CrB (orange line). The blue line is the total muonic neutrino flux from the RS Oph outburst. The IceCube’s upper limit on muon neutrino flux is green line (Pizzuto et al., 2021).
4 Conclusions
Using MESA, we simulate the evolution for nova outburst with CO WD mass ranging from 0.6 to 1.1 $\rm M_{\odot}$, and 1.1 $\rm M_{\odot}$ ONe WD. We calculate the low-energy nuclear and thermal neutrino luminosity, and give neutrino H-R diagram as well as low-energy nuclear and thermal neutrino luminosity curves. Our results show: During the accretion phase, the nuclear neutrino is mainly produced by the pp chains, and the thermal neutrino is mainly dominated by plasma decay. Neutrinos are produced mainly by thermal processes. Compared with photons, neutrinos carry away less energy. However, during the TNR, the temperature of the nuclear reaction zone on the surface of the nova exceeds $7\times 10^{7}\rm\ K$, the nuclear neutrino is mainly dominated by the CNO cycle, and the thermal neutrino is mainly dominated by photon neutrino, nuclear neutrino will far more exceed thermal neutrino. Meanwhile, photon is easily trapped during TNR, the low-energy nuclear and thermal neutrino luminosity rises rapidly and is much larger than the photon luminosity.
The $M_{\mathrm{WD}}$, $\dot{M}$, $T_{\mathrm{C}}$, and mixing have significant effects on the low-energy nuclear neutrino luminosity of nova. The temperature of the nuclear reaction zone on the surface of a WD increases with mass, so does the low-energy nuclear neutrino luminosity. The smaller the $\dot{M}$, the longer the outburst interval, the more accreted material, and the more intense the outburst, so the greater the low-energy nuclear neutrino luminosity. The model that does not take into account mixing has a higher low-energy nuclear neutrino luminosity because it has a denser accretion layer and a higher temperature. Cold nova model has lower low-energy neutrino luminosity in the accretion phase and higher low-energy nuclear neutrino luminosity during the TNR. It is because its core and shell are both cooler during the accretion phase. And it takes longer to reach the TNR, so more material is accreted, the shell is hotter, and the eruptions are more intense.
We predict that the low-energy nuclear neutrino peak luminosity of the next outburst of the T CrB will be $2.7\times 10^{8}\ L_{\nu,\odot}$ and a low-energy nuclear neutrino outburst duration of 88 days. We also predict the low-energy nuclear and high-energy hadronic neutrino flux that will reach Earth during the next outburst of T CrB. The results show that it is unlikely for current-generation Icecube to observe the neutrino flux produced from the hadronic processes during the T CrB next outburst.
Acknowledgements
This work received the generous support of the National Natural Science Foundation of China under grants No.12163005, U2031204, 12373038, 12288102 and 12263006, the science research grants from the China Manned Space Project with No. CMSCSST-2021-A10, and the Natural Science Foundation of Xinjiang No. 2022D01D85 and 2022TSYCLJ0006, and the Major Science and Technology Program of Xinjiang Uygur Autonomous Region under grant No.2022A03013-3.
Data Availability
The data generated, analysed and presented in this study are available from the corresponding authors on reasonable request.